# FoC: Figure out the Cryptographic Functions in Stripped Binaries with LLMs

### v2

@@ 1. Introduction
Cryptography-Oriented Heuristics Methods. There are some methods specifically designed to provide semantic information about cryptographic algorithms present in binary code. Current methods for cryptographic algorithm identification, which provide the primitive classes contained within binaries, utilize various approaches such as constants (findcrypt_yara, ; IDA_Signsrch, ; ReFormat_2009_ESORICS, ), statistics (grobert_2011_RAID, ; grap_GreHack_2017, ; KHUNT_CCS_2018, ; Kochberger_ICSSA_2018, ), structures (RuoxuZhao_Inscrypt_2014, ; JizhongLi_2014_ComEngDes, ; Lestringant_AsiaCCS_2015, ; Aligot_CCS_2012, ; CryptoHunt_SP_2017, ; wherescrypto_USENIX_2021, ), and others. These methods employ static or dynamic analysis to identify distinct patterns to detect cryptographic implementations within binaries. However, various factors affecting binary code (e.g., hidden constant features and compilation optimizations) can undermine the effectiveness of these methods. Additionally, the semantic information provided by these primitive classes is also quite limited.
@@
Binary Code Similarity Detection. Binary Code Similarity Detection (BCSD) is also a potential technical route. These methods (CoP_TSE_2017, ; BCSD_SAFE_2019, ; BCSD_DEEPBINDIFF, ; BCSD_PalmTree_CCS_2021, ; BCSD_TREX_TSE_2023, ; BCSD_jTrans_ISSTA2022, ; yang2023asteria, ) typically generate embeddings of binary functions to capture code features, thereby measuring the similarity between two binary functions. This allows us to identify functions in a known cryptographic function library that are similar to an unknown cryptographic function. Additionally, in contrast, the prediction of generative language models inherently struggles to reflect minor changes in binary code. However, distinguishing between two similar code is an essential ability, especially to recognize patched and vulnerable cryptographic functions. The function embeddings generated by the BCSD method can highly sensitively reflect any code changes, thus, to some extent, compensating for the shortcomings of generative models. Although current BCSD methods have demonstrated promising results on general datasets, limited attention has been given to the domain of cryptographic binary.
@@ 2.2. Challenges
C1: Available Cryptographic Binary Dataset. One of the fundamental challenges is the lack of publicly available comprehensive datasets tailored for the analysis of cryptographic functions in stripped binaries. There are dozens of cryptographic algorithms currently in the public domain, and a complete collection of their implementations is challenging. Moreover, since we focus on the binary domain, compiling the collected source code into binary code is an essential and labor-intensive task. In short, the absence of publicly available datasets hinders research on the current issue.
C2: Variety of Cryptographic Implementation. Mainstream cryptographic algorithms, such as ECC (ECC_adalier2015efficient, ) and AES (AES_pub2001197, ), serve as the foundation of computer security and must adhere to stringent specifications and standards to ensure both security and interoperability. However, practical implementations can differ significantly due to variations in the mathematical kernel and work mode employed.
@@
Therefore, it is crucial to collect as many implementations of cryptographic algorithms as possible. These implementations may be based on different standards and protocols, and designed for various platforms and purposes. Additionally, differences can arise even within the same algorithm specification due to developers’ programming styles. For instance, developers might decompose a cryptographic algorithm with complex operations into multiple functions or, conversely, merge several simple operations into a single function. Such variations can undermine the effectiveness of certain control-flow or data-flow analysis methods. Furthermore, since we are dealing with binary code, it is essential to account for differences introduced by compilation environments, including optimizations (e.g., loop unrolling, function inlining), and target architectures. Thus, even identical source code can yield vastly different binary representations.
@@
C3: Well-built Semantic Labels. For effective training and evaluation, it is necessary to have high-quality semantic labels that accurately describe the functionality of cryptographic functions in binaries. However, the creation of these labels poses significant challenges. First, mapping the semantic information present in the source code to the binary code is a feasible solution, but the function names and primitive classes in the source code are too brief to carry enough semantic details. Source code comments can provide valuable insights in many cases, but in real-world projects, comments are often missing or incomplete. Additionally, for large datasets containing millions of binary functions, manually annotating them is time-consuming and requires a high degree of expertise, which is impractical. Therefore, an automated method to generate accurate semantic labels is crucial.
@@
C4: Cross-version Awareness. Given the central role of cryptographic libraries in computer security and their widespread use, any vulnerability in their implementation can lead to unacceptable damage, such as the Heartbleed vulnerability (CVE-2014-0160, ). As a result, cryptographic algorithms are frequently updated to address vulnerabilities or improve performance, and such updates may be tiny code changes, such as applying a single-line patch to fix a critical vulnerability. For any analysis method, being sensitive to these tiny code changes is crucial, and achieving such cross-version awareness is particularly challenging in the context of purely generative models.
@@ 2.3. Existing Techniques
Technique Method C1 C2 C3 C4 Technique Method C1 C2 C3 C4
@@
Cryptography-Oriented Heuristics Methods. Cryptographic algorithm identification methods have been a focal point of research for over two decades (ChenxiaZhao_JPCS_2021, ). Numerous methods based on program structure have been proposed, especially the data-flow graph (DFG). For instance, Aligot (Aligot_CCS_2012, ) identifies the data-flow loops within the execution trace, while CryptoHunt (CryptoHunt_SP_2017, ) and Wherescrypto (wherescrypto_USENIX_2021, ) construct DFGs with symbolic execution. These methods rely on manually designed graph patterns with known implementations, making them incapable of overcoming C2. It means that these methods are not robust to any factors that make the binary code change.
Certain cryptographic algorithms contain noticeable features, such as constants (e.g., S-box) and statistical attributes. FindCrypt2 (FindCrypt2, ), Signsrch (IDA_Signsrch, ), and findcrypt-yara (findcrypt_yara, ) are three popular tools that identify the cryptographic algorithms present in binary files based on constant features. However, these methods will fail facing intentionally altered implementations or algorithms where constant values do not exist. Coarse-grained results at the file level (i.e., algorithm classes) fail on C3. Advanced methods like CryptoKnight (CryptoKnight_Information_2018, ) and GENDA (GENDA_JISA_2022, ) use CNN and GNN, respectively, to learn function semantics to predict primitive classes. Nonetheless, CryptoKnight uses a dataset almost exclusively from OpenSSL (OpenSSL, ), and GENDA’s dataset is sourced from only four cryptographic algorithm libraries. These limitations in data diversity contribute to their inadequacies in addressing C2.
@@
Binary Code Similarity Detection. BCSD methods aim to compare the degree of similarity between two binary code snippets, and have the potential to overcome C4, i.e., cross-version awareness. BCSD plays a vital role in various security tasks, including malware detection, plagiarism detection, and patching analysis. By identifying similar or identical code in different binaries, the BCSD approach enables security analysts to track code reuse and identify potential vulnerabilities.
@@ 2.4. Large Language Model and Our Motivation
Recently, Large Language Models (LLMs) have captured significant attention from both academia and industry due to their remarkable capabilities. Typically, LLMs refer to language models that contain tens of billions or more parameters (LLMsSurvey_2023_arxiv, ), trained on vast amounts of textual data using extensive computational resources. Notable examples include GPT-3 (GPT3_2020_NIPS, ), PaLM (PaLM_2022_arxiv, ), and LLaMA (LLaMA_2023_arxiv, ), which have shown impressive performance across a range of natural language processing tasks. Under this research boom, LLMs specifically focused on programming languages have also been proposed rapidly, including Codex (Codex_2021_arxiv, ), GPT-J (gpt-j-6b_model_2021, ), GPT-NeoX (GPT-NeoX-20B_ACL_2022, ), CodeT5+ (CodeT5P_2023_arxiv, ) PolyCoder (PolyCoder_PLDI_2022, ), WizardCoder (WizardCoder_2023_arxiv, ), CodeLlama (codellama2023rozière, ), and so on. These models have demonstrated exceptional code comprehension capabilities.
Given the demonstrated powerful understanding capabilities of LLMs in natural language (NL) and programming language (PL) tasks, it has become possible to devise an automated pipeline to create high-quality semantic descriptions for the cryptographic function in source code. This approach mitigates the challenge of creating semantic labels (C3). Consequently, we can confidently gather a rich cryptographic binary dataset to tackle challenges C1 and C2. Additionally, a recent evaluation (Binary_Code_Summarization_Benchmarking, ) highlights that LLMs perform significantly worse on summarizing binary code compared to source code. Therefore, using the dataset, we propose developing a binary-specific LLM that generates comprehensive descriptions of cryptographic functions in stripped binaries, which potentially addresses C3. To mitigate the generative model’s limitations regarding C4, we further incorporate a BCSD module based on our binary LLM to create the cross-version aware embedding representation.
@@ 3.1. Cryptographic Dataset Collection
Figure 1. (Left) Development of open-source cryptography repositories we investigated. The train data highlighted in blue, the test data highlighted in red. (Right) The code overlap between cryptography repositories.
As discussed in C1 and C2 in Section 2.2, the lack of publicly available and implementation-diverse cryptographic binary datasets has prevented research in the current field. Therefore, we here collect a comprehensive cryptographic binary dataset for building our methods and stimulating further research. Our focus is on studying widely-used cryptographic algorithm libraries, popular cryptographic implementations, and cryptographic modules within large projects, all of which are written in the C programming language.
@@
As shown in Figure 2, for each project in the dataset, we employ cross-compilation to obtain their binaries under different compilation environments. Specifically, we employ two compilers, GCC-11.2.0 and Clang-13.0, with four different optimization options, i.e., O0-O3. These projects are compiled for six different target architectures, including x86_32, x86_64, arm_32, arm_64, mips_32, and mips_64. Subsequently, we strip the binaries to make them consistent with release versions in real-world scenarios. IDA Pro (IDA_PRO, ) is used to decompile the binary files. We perform deduplication on all functions in our dataset, using the MD5 digest of binary functions, to avoid data redundancy. Our dataset contains considerable cryptographic algorithms and takes account of the complex compilation environments in the real world, which allows us to overcome C1 and C2. The statistical information of the dataset is shown in Table 2.
@@ 3.2. Automated Semantic Labels Creation
As we described in C3, creating NL semantic labels for extensive binary functions is a challenging issue. As shown in Figure 2, we can establish correspondences between binary code and source code using DWARF (Dwarfformat, ) debugging information and offset addresses in the binary file, but comprehending the source code itself is not straightforward.
@@
To mitigate this problem, inspired by recent research works (zhang2023llmaaa, ; xiao2023freeal, ) that utilize LLMs to perform data annotation tasks with reasonable reliability, we leverage ChatGPT (chatgpt-GPT3.5-2022, ), an advanced general LLM, to automatically generate summaries as comments, outlining the function’s purpose and functionality. Specifically, we use the metadata extracted from the source code for each function to build the prompt shown in Figure 3. We prompt the LLM to generate only one-sentence summaries to try to avoid including too much information that is not present in the pseudo-code, such as variable names and macro definitions, which potentially bias the model. On the other hand, a one-sentence summary is easier to read and understand.
@@ 3.3. Keyword-based Discriminator
It is essential to assess whether these summaries align with the facts. Therefore, we propose a keyword-based discriminator to judge the correctness of generated summaries on the crucial cryptographic semantics. Specifically, as shown in Table 4, we have defined three categories and specific classes. We use whole-word matching to retrieve class-related keywords. However, a class is often written in multiple forms in summary or source code. For example, ’3des’ might be written as ’triple-des’, ’triple des’, ’tripledes’, ’desede’, and ’des-ede’ etc. To get the most accurate results possible, we perform extensive manual inspections to comprehensively summarize the different forms that a class may exist in, as shown in Table 5. The discriminator considers these different forms as the same class, and it passes the inspection only when it obtains the same class between the generated summary and the source code.
The results indicate that more than 85% of the generated summary passed and they were retained in the end. We further evaluate the textual consistency between developer-written comments and model-generated summaries. The results show that they have a 43.55% ROUGE-L score, which means a high degree of consistency. And Table 3 shows a few generated labels. Overall, for each function in our dataset, we obtain its binary code, source code, and semantic labels, which highly align with the facts. In this way, we address the challenges presented in C3.

### v3

@@ 1. Introduction
Cryptography-Oriented Heuristics Methods. There are some methods specifically designed to provide semantic information about cryptographic algorithms present in binary code. Current methods for cryptographic algorithm identification, which provide the primitive classes contained within binaries, utilize various approaches such as constants (findcrypt_yara, ; IDA_Signsrch, ; ReFormat_2009_ESORICS, ), statistics (grobert_2011_RAID, ; grap_GreHack_2017, ; KHUNT_CCS_2018, ; Kochberger_ICSSA_2018, ), structures (RuoxuZhao_Inscrypt_2014, ; JizhongLi_2014_ComEngDes, ; Lestringant_AsiaCCS_2015, ; Aligot_CCS_2012, ; CryptoHunt_SP_2017, ; wherescrypto_USENIX_2021, ), and others. These methods employ static or dynamic analysis to identify distinct patterns to detect cryptographic implementations within binaries. However, various factors affecting binary code (e.g., hidden constant features and compilation optimizations) can undermine the effectiveness of these methods. Additionally, these methods can usually only identify simple primitive classes, and lack a deep understanding of the complex binary code structure, providing very limited semantic information to human analysts.
@@
Binary Code Similarity Detection. Binary Code Similarity Detection (BCSD) is also a potential technical route. These methods (CoP_TSE_2017, ; BCSD_SAFE_2019, ; BCSD_DEEPBINDIFF, ; BCSD_PalmTree_CCS_2021, ; BCSD_TREX_TSE_2023, ; BCSD_jTrans_ISSTA2022, ; yang2023asteria, ) typically generate embeddings of binary functions to capture code features, thereby measuring the similarity between two binary functions. This allows us to identify functions in a known cryptographic function library that are similar to an unknown cryptographic function. Additionally, in contrast, the prediction of generative language models inherently struggles to reflect minor changes in binary code. However, distinguishing between two similar code is an essential ability, especially to recognize patched and vulnerable cryptographic functions. The function embeddings generated by the BCSD method can highly sensitively reflect any code changes, thus, to some extent, compensating for the shortcomings of generative models. Although current BCSD methods have demonstrated promising results on general datasets, limited attention has been given to the domain of cryptographic binary. In addition, the effectiveness of the BCSD methods is limited by the scope and comprehensiveness of the known cryptographic function library, and performs poorly when dealing with unknown or variant binary functions.
@@ 2.2. Challenges
C1: Available and Diverse Cryptographic Binary Dataset. One of the fundamental challenges is the lack of publicly available comprehensive datasets tailored for the analysis of cryptographic functions in stripped binaries. There are dozens of cryptographic algorithms currently in the public domain, and a complete collection of their implementations is challenging. Moreover, since we focus on the binary domain, compiling the collected source code into binary code is an essential and labor-intensive task. In short, the absence of publicly available datasets hinders research on the current issue.
@@
In addition, mainstream cryptographic algorithms, such as ECC (ECC_adalier2015efficient, ) and AES (AES_pub2001197, ), serve as the foundation of computer security and must adhere to stringent specifications and standards to ensure both security and interoperability. However, practical implementations can differ significantly due to variations in the mathematical kernel and work mode employed. To advance research, it is crucial to collect as diverse implementations of cryptographic algorithms as possible. These implementations may be based on different standards and protocols, and designed for various platforms and purposes. Additionally, differences can arise even within the same algorithm specification due to developers’ programming styles. For instance, developers might decompose a cryptographic algorithm with complex operations into multiple functions or, conversely, merge several simple operations into a single function. Such variations can undermine the effectiveness of certain control-flow or data-flow analysis methods. Furthermore, since we are dealing with binary code, it is essential to account for differences introduced by compilation environments, including optimizations (e.g., loop unrolling, function inlining), and target architectures, which can lead to vastly different binary representations from identical source code.
@@
C2: Well-built Semantic Labels. For effective training and evaluation, it is necessary to have high-quality semantic labels that accurately describe the functionality of cryptographic functions in binaries. However, the creation of these labels poses significant challenges. First, mapping the semantic information present in the source code to the binary code is a feasible solution, but the function names and primitive classes in the source code are too brief to carry enough semantic details. Source code comments can provide valuable insights in many cases, but in real-world projects, comments are often missing or incomplete. Additionally, for large datasets containing millions of binary functions, manually annotating them is time-consuming and requires a high degree of expertise, which is impractical. Therefore, an automated method to generate accurate semantic labels is crucial.
@@
C3: Cross-version Awareness. Given the central role of cryptographic libraries in computer security and their widespread use, any vulnerability in their implementation can lead to unacceptable damage, such as the Heartbleed vulnerability (CVE-2014-0160, ). As a result, cryptographic algorithms are frequently updated to address vulnerabilities or improve performance, and such updates may be tiny code changes, such as applying a single-line patch to fix a critical vulnerability. For any analysis method, being sensitive to these tiny code changes is crucial, and achieving such cross-version awareness is particularly challenging in the context of purely generative models.
@@ 2.3. Existing Techniques
Technique Method C1 C2 C3 Technique Method C1 C2 C3
@@
  • ”✓” indicates that the method effectively addresses the corresponding challenge, ”✗” indicates that it does not, and ”✓” indicates that it partially addresses the challenge.
@@
Cryptography-Oriented Heuristics Methods. Cryptographic algorithm identification methods have been a focal point of research for over two decades (ChenxiaZhao_JPCS_2021, ). Numerous methods based on program structure have been proposed, especially the data-flow graph (DFG). For instance, Aligot (Aligot_CCS_2012, ) identifies the data-flow loops within the execution trace, while CryptoHunt (CryptoHunt_SP_2017, ) and Wherescrypto (wherescrypto_USENIX_2021, ) construct DFGs with symbolic execution. These methods rely on manually designed graph patterns with known implementations, making them incapable of overcoming C1. It means that these methods are not robust to any factors that make the binary code change.
Certain cryptographic algorithms contain noticeable features, such as constants (e.g., S-box) and statistical attributes. FindCrypt2 (FindCrypt2, ), Signsrch (IDA_Signsrch, ), and findcrypt-yara (findcrypt_yara, ) are three popular tools that identify the cryptographic algorithms present in binary files based on constant features. However, these methods will fail facing intentionally altered implementations or algorithms where constant values do not exist. Coarse-grained results at the file level (i.e., algorithm classes) fail on C2. Advanced methods like CryptoKnight (CryptoKnight_Information_2018, ) and GENDA (GENDA_JISA_2022, ) use CNN and GNN, respectively, to learn function semantics to predict primitive classes. Nonetheless, CryptoKnight uses a dataset almost exclusively from OpenSSL (OpenSSL, ), and GENDA’s dataset is sourced from only four cryptographic algorithm libraries. These limitations in data diversity contribute to their inadequacies in addressing C1.
@@
Binary Code Similarity Detection. BCSD methods aim to compare the degree of similarity between two binary code snippets, and have the potential to overcome C3, i.e., cross-version awareness. BCSD plays a vital role in various security tasks, including malware detection, plagiarism detection, and patching analysis. By identifying similar or identical code in different binaries, the BCSD approach enables security analysts to track code reuse and identify potential vulnerabilities.
@@ 2.4. Large Language Model and Our Motivation
Recently, Large Language Models (LLMs) have captured significant attention from both academia and industry due to their remarkable capabilities. Typically, LLMs refer to language models that contain tens of billions or more parameters (LLMsSurvey_2023_arxiv, ), trained on vast amounts of textual data using extensive computational resources. Notable examples include GPT-3 (GPT3_2020_NIPS, ), PaLM (PaLM_2022_arxiv, ), and LLaMA (LLaMA_2023_arxiv, ), which have shown impressive performance across a range of natural language processing tasks. Under this research boom, LLMs specifically focused on programming languages have also been proposed rapidly, including Codex (Codex_2021_arxiv, ), GPT-J (gpt-j-6b_model_2021, ), GPT-NeoX (GPT-NeoX-20B_ACL_2022, ), CodeT5+ (CodeT5P_2023_arxiv, ), PolyCoder (PolyCoder_PLDI_2022, ), WizardCoder (WizardCoder_2023_arxiv, ), CodeLlama (codellama2023rozière, ), and so on. These models have demonstrated exceptional code comprehension capabilities.
Additionally, for stripped binaries, semantic information such as types, variable names, function names, and even structural details like instruction and function boundaries are discarded during the compilation and stripping process. This makes semantic analysis of binary code extremely challenging, as it involves recovering missing information, which is essentially creating something from nothing. In this context, traditional analysis techniques and small-scale models are limited in their effectiveness. LLMs, however, present a viable solution. Applications such as text-to-image (qin2024diffusiongpt, ) and text-to-video (huang2024free, ) generation have shown that LLMs can compensate for and restore missing information, which is especially valuable for analyzing stripped binaries with significant information gaps. Although traditional analysis techniques have reached a bottleneck, LLMs are rapidly advancing, offering great potential for further improvement.
In summary, given the demonstrated powerful understanding capabilities of LLMs in natural language (NL) and programming language (PL) tasks, it has become possible to devise an automated pipeline to create high-quality semantic descriptions for the cryptographic function in source code. This approach mitigates the challenge of creating semantic labels (C2). Consequently, we can confidently gather a rich cryptographic binary dataset to tackle challenges C1. Additionally, a recent evaluation (Binary_Code_Summarization_Benchmarking, ) highlights that LLMs perform significantly worse on summarizing binary code compared to source code. Therefore, using the dataset, we propose developing a binary-specific LLM that generates comprehensive descriptions of cryptographic functions in stripped binaries, which potentially addresses C2. To mitigate the generative model’s limitations regarding C3, we further incorporate a BCSD module based on our binary LLM to create the cross-version aware embedding representation.
@@ 3.1. Cryptographic Dataset Collection
Figure 1. (Left) Development of open-source cryptography repositories we investigated. The train data highlighted in blue, the test data highlighted in red. (Right) The code overlap between cryptography repositories, where the numbers represent the number of functions that are shared or unique among the repositories.
As discussed in C1 in Section 2.2, the lack of publicly available and implementation-diverse cryptographic binary datasets has prevented research in the current field. Therefore, we here collect a comprehensive cryptographic binary dataset for building our methods and stimulating further research. Our focus is on studying widely-used cryptographic algorithm libraries, popular cryptographic implementations, and cryptographic modules within large projects, all of which are written in the C programming language.
@@
As shown in Figure 2, for each project in the dataset, we employ cross-compilation to obtain their binaries under different compilation environments. Specifically, we employ two compilers, GCC-11.2.0 and Clang-13.0, with four different optimization options, i.e., O0-O3. These projects are compiled for six different target architectures, including x86_32, x86_64, arm_32, arm_64, mips_32, and mips_64. Subsequently, we strip the binaries to make them consistent with release versions in real-world scenarios. IDA Pro (IDA_PRO, ) is used to decompile the binary files. We perform deduplication on all functions in our dataset, using the MD5 digest of binary functions, to avoid data redundancy. Our dataset contains considerable cryptographic algorithms and takes account of the complex compilation environments in the real world, which allows us to overcome C1. The statistical information of the dataset is shown in Table 2.
@@ 3.2. Automated Semantic Labels Creation
As we described in C2, creating NL semantic labels for extensive binary functions is a challenging issue. As shown in Figure 2, we can establish correspondences between binary code and source code using DWARF (Dwarfformat, ) debugging information and offset addresses in the binary file, but comprehending the source code itself is not straightforward.
@@
To mitigate this problem, inspired by recent research works (zhang2023llmaaa, ; xiao2023freeal, ) that utilize LLMs to perform data annotation tasks with reasonable reliability, we leRverage ChatGPT (chatgpt-GPT3.5-2022, ), an advanced general LLM, to automatically generate summaries as comments, outlining the function’s purpose and functionality. Specifically, we use the metadata extracted from the source code for each function to build the prompt shown in Figure 3. We prompt the LLM to generate only one-sentence summaries to try to avoid including too much information that is not present in the pseudo-code, such as variable names and macro definitions, which potentially bias the model. On the other hand, a one-sentence summary is easier to read and understand.
@@ 3.3. Keyword-based Discriminator
It is essential to assess whether these summaries align with the facts. Therefore, we propose a keyword-based discriminator to judge the correctness of generated summaries on the crucial cryptographic semantics. Specifically, as shown in Table 4, we have defined three categories and specific classes. We use whole-word matching to retrieve class-related keywords. However, a class is often written in multiple forms in summary or source code. For example, ‘3des’ might be written as ‘triple-des’, ‘triple des’, ‘tripledes’, ‘desede’, and ‘des-ede’ etc. To get the most accurate results possible, we perform extensive manual inspections to comprehensively summarize the different forms that a class may exist in, as shown in Table 5. The discriminator considers these different forms as the same class, and it passes the inspection only when it obtains the same class between the generated summary and the source code.
The results indicate that more than 85% of the generated summary passed and they were retained in the end. We further evaluate the textual consistency between developer-written comments and model-generated summaries. The results show that they have a 43.55% ROUGE-L score, which means a high degree of consistency. And Table 3 shows a few generated labels. Overall, for each function in our dataset, we obtain its binary code, source code, and semantic labels, which highly align with the facts. In this way, we address the challenges presented in C2.